OpenStudy (anonymous):

Factor: x2 - 9x + 12 A. (x + 1)(x - 12) B. (x + 2)(x + 6) C. (x - 2)(x - 6) D. prime

OpenStudy (callisto):

x2 - 9x + 12 <-correct question?

OpenStudy (anonymous):

yupp

OpenStudy (anonymous):

On multiple choice questions, a good strategy is to solve each answer to see what fits

OpenStudy (anonymous):

So if you can't easily factor the question, multiply the answers out!

OpenStudy (callisto):

It can't be factorised then. At least the factored form is not in the above choices

OpenStudy (callisto):

@toobulkeh Good point!

OpenStudy (anonymous):

so C?

OpenStudy (anonymous):

C would be something like: \[x ^{2} -8x + 12\]

OpenStudy (anonymous):

Right option mot mentioned

OpenStudy (anonymous):

*not

OpenStudy (anonymous):

b!

OpenStudy (anonymous):

Multiply B out, what is the polynomial that you get?

OpenStudy (callisto):

A. (x + 1)(x - 12) = x^2 -12x+x -12 = x^2-11x-12 B . (x + 2)(x + 6) = x^2 +2x+6x+12 = x^2 +8x+12 C. (x - 2)(x - 6) = x^2 -2x-6x+12 = x^2 -8x +12